# Not All LLM-Generated Data Are Equal: Rethinking data weighting in Text Classification

### v1

@@ G.3 Analysis of Checker Scores and Weights
Figure 4: Average Quality Checker Score, Diversity Checker Score, and Weights of IMP-Loss for Financial Dataset: Comparison between Original, Swapped Label, and Duplicated Data.
Figure 5: Average Quality Checker Score, Diversity Checker Score, and Weights of IMP-Loss for Tweet Irony Dataset: Comparison between Original, Swapped Label, and Duplicated Data.
@@
The Quality Checker Score (blue bar) is highest for the original data and significantly lower for the swapped label data, indicating that the model correctly identifies the labels as lower quality. The Diversity Checker Score (where lower values are better) is lower for the original data than the duplicated data, indicating the impact of duplication on diversity. Additionally, according to the diversity checker, the swapped label data exhibits the highest diversity. This is because the data points with changed labels are significantly different from other data in the dataset, thus enhancing their diversity score. Consequently, the weights assigned to the original data are higher compared to the swapped label data and the duplicated data, demonstrating the effectiveness of IMP-Loss in recognizing and appropriately weighting high-quality, diverse data.
@@
Figure 6: Average Diversity Checker Score of DIMP-Loss for Original, Swapped label, and Duplicated Data on the Financial Dataset across Epoch.
Figure 7: DIMP-Loss Weight for Original, Swapped label, and Duplicated Data on the Financial Dataset across Epoch.
In contrast, Figure 6 shows the diversity scores $\hat{P}(y|\mathbf{x};\theta_{t})$ and the weights for the DIMP-Loss method on Financial benchmark across epoch, where the diversity checker is the training model itself. The Diversity Checker Score (lower is better) is also lower for the original data than the duplicated data. In the end, The weights (Figure 7) assigned by the DIMP-Loss method are consistently higher for the original data than the swapped label and duplicated data across epochs. This pattern aligns with the results observed for the IMP-Loss method.

### v2

@@ G.3 Analysis of Checker Scores and Weights
The Quality Checker Score is highest for the original data and significantly lower for the swapped label data, indicating that the model correctly identifies the labels as lower quality. The Diversity Checker Score (where lower values are better) is lower for the original data than the duplicated data, indicating the impact of duplication on diversity. Additionally, the swapped label data achieves the highest diversity because the altered labels create data points that are substantially distinct from the rest of the dataset. Similarly, the unrelated input data exhibits relatively high diversity due to its out-of-domain nature. However, data points from both the swapped label and unrelated input categories have low Quality Checker Scores, resulting in their lower assigned weights. Consequently, the weights assigned to the original data are higher compared to the swapped label data and the duplicated data, demonstrating the effectiveness of IMP-Loss in recognizing and appropriately weighting high-quality, diverse data.
Figure 4: Average Quality Checker Score, Diversity Checker Score, and Weights of IMP-Loss for Financial Dataset: Comparison between Original, Swapped Label, Duplicated Data and Unrelated Input Data.
Figure 5: Average Quality Checker Score, Diversity Checker Score, and Weights of IMP-Loss for Tweet Irony Dataset: Comparison between Original, Swapped Label, and Duplicated Data
@@
Figure 6: Average Diversity Checker Score of DIMP-Loss for Original, Swapped label, Unrelated input data, and Duplicated data on the Financial Dataset across Epoch.
Figure 7: DIMP-Loss Weight for Original, Swapped label, Unrelated input data, and Duplicated Data on the Financial Dataset across Epoch.
In contrast, Figure 6 shows the diversity scores $\hat{P}(y|\mathbf{x};\theta_{t})$ and the weights for the DIMP-Loss method on Financial benchmark across epoch, where the diversity checker is the training model itself. The Diversity Checker Score (lower is better) is also lower for the original data than the duplicated data and Unrelated input data. In the end, The weights (Figure 7) assigned by the DIMP-Loss method are consistently higher for the original data than the swapped label, Unrelated input data, and duplicated data across epochs. This pattern aligns with the results observed for the IMP-Loss method.